Circular No. 5822

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

HD 173637 A. J. Landolt, Louisiana State University, writes: “The star HD 173637

has been classified as a B0e star with an ‘almost continuous spectrum’ (Neubauer 1943, Ap.J. 97, 300) and as B1 IV (Morgan et al. 1955, Ap.J. Suppl. 2, 41). Insofar as is known, the star has been constant over many years, and has been used successfully as a photometric standard star (Landolt 1983, A.J. 88, 439). HD 173637 was found in an apparent outburst of > 0.4 mag on June 25.27 UT with the Yale 1.0-m telescope at Cerro Tololo Interamerican Observatory. Rough quick-look magnitudes and color indices are: V = 8.96, B–V = +0.37, U–B = –0.8.”

SUPERNOVA 1993J IN NGC 3031 R. Martin writes: “The data taken of SN 1993J with the Isaac New-

ton Group of Telescopes (INGT) and Carlsberg Automatic Meridian Circle of the Observatorio del Roque de los Muchachos, are being reduced at the INGT on La Palma and at the Royal Greenwich Observatory in Cambridge. These reduced data are being written into a public-domain archive that is available via anonymous ftp over Internet. The site name for access is ftp.ast.cam.ac.uk and the data are organised in the directories that are in the tree /sn1993j/reduced/.... The file README gives information on the data, instruments, telescopes, etc. The file /sn1993j/reduced/acknowledgements.doc gives the acknowledgement that should be used when publishing articles based on this archived data. Further information on the archive can be obtained from Ralph Martin or Nic Walton, Royal Greenwich Observatory, Madingley Road, Cambridge CB3 0EZ, England.”

R. Pressberger and H. M. Maitzen, Institute for Astronomy, Vienna, report further Strömgren photometry (cf. IAUC 5816), given as before in the sense (SN 1993J – GSC 4383.0928): June 1.89 UT, u = +2.1, v = +1.44, b = +1.25, g_{1} = +0.97, g_{2} = +0.83, y = +0.61.

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Total V magnitude estimates (cf. IAUC 5798) from H. Mikuž, Ljubl-

jana, Slovenia (0.20-m Baker-Schmidt camera + ST-6 CCD): May 22.89 UT, 14.1; June 18.87, 14.1.

1993 June 25 Daniel W. E. Green